OpenStudy (anonymous):

what is x when 3sinx-4sinx=2

OpenStudy (nincompoop):

think of 3x - 4x = 2 -x = 2 right? similarly -sin(x) = 2

OpenStudy (anonymous):

I miswritten it, what I meant was 3sinx-4cosx=2

OpenStudy (anonymous):

how do I express 4cosx in terms of sinx or should I express 3sinx in terms of cosx?

OpenStudy (anonymous):

What identity?